# UrbanWorld: An Urban World Model for 3D City Generation

### v1

@@ 3.4 MLLM-assisted urban scene refinement
After urban asset rendering, UrbanWorld automatically reorganizes the assets guided by the location information extracted from the real OSM data, effectively recovering the original urban layout. Inspired by the standard operation process of human designing, where experts will take an overview of the scene and make minor adjustments. To mimic such effort, UrbanWorld activates urban MLLM again to scrutinize the crafted 3D urban scenes and texture details. We prompt the MLLM to identify inconsistencies between the scene imagery and previous design prompts and examine whether the texture is as realistic as in the real world. Finally, the urban MLLM will provide sophisticated suggestions for further refinement, including elements to be modified and refined design prompts. Then the rendering module will be activated and the involved elements will be rendered under the refined text prompts and updated in the scene. With such a refinement process, UrbanWorld can further align the generated urban environment to the real cities. We provide a visualization example showing the evolution of the created urban environments in Figure 3. It can be seen that UrbanWorld works in an iterative refinement manner to create high-fidelity urban environments, where the low-quality textures will be automatically identified and refined with the powerful urban MLLM.
@@ 4 Experiments
Figure 4: Qualitative comparisons of generated 3D urban environments from Infinicity, CityGen (results from the original paper since the code is not open-source), CityDreamer and UrbanWorld. By comparison, our method can craft more diverse 3D urban scenes with high-fidelity textures following user instructions.
In this section, we first introduce the experimental setup and implementation details (see Section 4.1), and then we present the generation results of UrbanWorld for qualitative estimation (see Section 4.2). Finally, we provide some quantitative evaluations of the created urban environments to demonstrate the superiority of UrbanWorld (see Section 4.3).
@@ 4.1 Implementation Details
UrbanWorld incorporates three key techniques: Blender as the professional 3D modeling software, diffusion-based rendering and MLLM-empowered scene design and refinement. Specifically, we use Blender-3.2.2 for Linux systems and the compatible Blosm addon to handle the OSM transformation. In terms of diffusion-based rendering, we utilize Stable Diffusion-1.5 (Rombach et al., 2022) as the fundamental diffusion backbone, together with ControlNet-Depth (Zhang et al., 2023) in the generation of multi-view images. We also introduce IP-Adapter (Ye et al., 2023) to support taking reference images as the additional generation condition. We use ControlNet-inpainting (Zhang et al., 2023) as the diffusion controller in the UV texture refinement and ContrlNet-tile (Zhang et al., 2023) in the realness enhancement stage. For the hyper-parameter settings in the diffusion-based rendering, we set the number of camera views $N=4$, which can basically satisfy the rendering needs of most urban assets. The number of inference steps in all diffusion processes is set as 30 by default. The UV maps of 3D assets are unwrapped in the “smart projection” mode operated in Blender. As for the urban MLLM, we fine-tune LLaVA-1.5 on a curated dataset with  100K streetview image-text data pairs. All experiments are conducted on a single NVIDIA Tesla A100 GPU.
4.2 Qualitative Results
We present some generation results of UrbanWorld in Figure 4, including various representative urban scenes, including educational places, commercial blocks and residential areas. For intuitive comparison, we also provide some generation samples from Infinicity (Lin et al., 2023), CityGen (Deng et al., 2023) and CityDreamer (Xie et al., 2024). The results of Infinicity and CityGen are taken from the original paper because the codes are not open-source. It can be seen that scenes from Infinicity are short of clear textures and well-maintained building structures. Scenes from CityGen feature homogeneous styles without clear characteristics of urban functions. Similarly, the visual appearance of urban elements (especially buildings) in the environments from CityDreamer lacks diversity and is hard to distinguish. Besides, there are also clear geometric distortions of the building boundaries in CityDreamer. These issues will pose great challenges for the real interactions between subjects and urban environments. For example, embodied agents are hard to be trained to conduct urban navigation because the surrounding elements are too similar to recognize. By comparison, the urban elements created by UrbanWorld possess distinct functional characteristics, benefiting from the high controllability with text and reference image prompts. Besides, the overall scene is more authentic and visually harmonious, demonstrating the effectiveness of intelligent scene planning and design of the MLLM.
@@ 4.3 Quantitative Evaluations
Method Depth Error (↓) Homogeneity Index (↓) Realistic Score (↑)
SceneDreamer (Chen et al., 2023) 0.152 0.746 5.6
CityDreamer (Xie et al., 2024) 0.147 0.745 6.0
UrbanWorld 0.089 0.683 6.8
To better demonstrate the superior generation performance of UrbanWorld, in this section, we provide quantitative results from the following three aspects:
Depth error. Depth error (DE) is utilized to evaluate the 3D geometry accuracy, following the implementation of EG3D (Chan et al., 2022) and CityDreamer (Xie et al., 2024). Specifically, we use the pre-trained depth estimation model (Ranftl et al., 2020) to obtain the “ground truth” of depth maps via density accumulation. DE is then calculated as the L2 distance between the normalized predicted depth maps and the “ground truth”. The final result is averaged on the result from 100 captured frames of generated urban scenes.
Homogeneity index. Realistic cities are featured by complex elements with diverse visual appearances, indicating various functional uses of different urban areas. In order to capture this key character, we propose to evaluate the homogeneity of generated scenes, mainly measuring the variance of different urban scenes. To be specific, we first extract the visual feature of each generated scene image with ResNet (He et al., 2016). The homogeneity index is then calculated as the averaged cosine similarity of each pair of scenes in the feature space. The smaller value of the homogeneity index means a higher diversity of generated urban environments.
Realistic score. Another important aspect to evaluate is the realness of generated urban environments, which is challenging even for human designers. Here we introduce an off-the-shelf MLLM (GPT-4o) as the evaluator to score the realistic level and quality of generated urban elements (ranging from 1 to 10). A higher realistic score indicates better realness and fidelity of generated textures of urban assets.
@@
We compare our model with some existing methods for urban scene generation, mainly including SceneDreamer (Deng et al., 2023) and Citydreamer Xie et al. (2024), representing the most advanced performance of automatic 3D urban environment generation. We don’t provide results from other methods such as CityGen (Deng et al., 2023), CityCraft (Deng et al., 2024) and SceneCraft (Hu et al., 2024) because the source codes are not open when this work was done.
@@
From the results presented in Table 2, it can be seen that UrbanWorld outperforms on each quantitative metric compared with baselines. Statistically, UrbanWorld achieves 39.5% on depth error compared with the most competitive baseline, demonstrating the geometry-preserving ability of UrbanWorld. By comparison, rendering methods such as SceneDreamer and CityDreamer can produce visually appealing urban scenes, but commonly lose geometry consistency. In terms of the homogeneity index, UrbanWorld has 8.3% improvement compared with baselines. Consistent with the observation on the qualitative results in Section 4.2, the generated scenes from existing methods exhibit high homogeneity, limited to the style of training data. By comparison, UrbanWorld can produce more diverse urban environments according to user instruction, effectively achieving customized creation. This makes it real for crafting any type of 3D urban environment adapting to the needs of different scenes. Lastly, the created urban environments from UrbanWorld show outstanding realness and high texture fidelity, increasing 11.8% realistic score compared with baselines. The overall impression of scenes crafted by SceneDreamer and CityDreamer is harmonious, however, the texture quality with a closer look is unsatisfying. Differently, UrbanWorld conducts element-wise texture rendering based on the specific 3D mesh, ensuring the geometric matching and fidelity of generated textures.
4.4 Ablation study
@@
Table 3: Study of the effectiveness of three key designs in UrbanWorld: MLLM-empowered urban scene design, texture enhancement and MLLM-assisted scene refinement.
Method Depth Error (↓) Homogeneity Index (↓) Realistic Score (↑)
UrbanWorld 0.089 0.683 6.8
w/o scene design 0.104 0.701 6.1
w/o texture enhancement 0.125 0.687 6.5
w/o scene refinement 0.096 0.690 6.4
@@
To further validate the effectiveness of key designs in UrbanWorld, we conduct ablation studies and show the results in Table 3. We explore the influence of three designs on the performance, including MLLM-empowered urban scene design, texture enhancement and MLLM-assisted scene refinement. The results indicate that all these techniques contribute to the final generation performance. Specifically, the scene design from the MLLM contributes most to the realness of generated urban environments, benefiting from the powerful knowledge understanding and reasoning ability of MLLMs. The texture completion and enhancement operation has the most notable effect on the estimated depth error because better texture fidelity can help with geometric perception. Besides, the final scene refinement process leads to a gain in all evaluation metrics, further promoting the whole performance.
@@ 5 Conclusion
We propose UrbanWorld, the first generative urban world model to create realistic, customized and interactive 3D urban environments with flexible control conditions in a fully automatic manner. Integrating the powerful urban scene understanding ability of urban MLLM and the controllable generation ability of diffusion models, UrbanWorld can effectively craft high-fidelity and diverse urban environments outperforming existing urban scene generation methods. More importantly, the created urban environments can provide realistic and high-fidelity information feedback, ensuring embodied agents interact with and learn from the richly detailed, realistic urban world. We are contributing UrbanWorld as an open-source tool to benefit broad research communities, including but not limited to AI agents and embodied intelligence. We believe this work can pave a new way to efficiently establish the 3D virtual urban environment, accelerating the development of AGI.
For future work, we will continue to enhance the current version of UrbanWorld and there are mainly three directions. Firstly, we will supplement more elements into urban environment creation, such as persons and vehicles, further improving the richness and realness of crafted environments. Secondly, we will conduct experiments on interactive tasks such as visual recognition and navigation for embodied agents to validate the real usability of created 3D environments. Lastly, we are organizing the codes and the generated 3D urban environment assets for release. We expect to make UrbanWorld an open-source toolkit for convenient use in various real applications.

### v2

@@ 3.4 MLLM-assisted Urban Scene Refinement
After urban asset rendering, UrbanWorld automatically reorganizes the assets guided by the location information extracted from the metadata of the urban layout, effectively recovering the original urban layout. Inspired by the standard operation process of human designing, where experts will take an overview of the result and conduct iterative adjustments. To mimic such an effort, UrbanWorld resorts to Urban MLLM again to scrutinize the crafted 3D urban environment, especially focusing on the texture details. Specifically, we prompt Urban MLLM to identify the inconsistencies between the generated result and the prompts for generation. Finally, Urban MLLM will provide sophisticated suggestions for further refinement, including elements to be modified and refined design prompts. Then the rendering module described in Section 3.3 will be activated and the involved elements will be rendered under the refined text prompts and updated in the 3D environment. With such a refinement process, UrbanWorld can further align the generated urban environment to user instructions. Here we provide a visualization example in Figure 3, showing the evolution of the created urban environments happened in the running process of UrbanWorld. It can be seen that UrbanWorld works in an iterative refinement manner to create high-fidelity urban environments, where the low-quality textures will be automatically identified and refined with the powerful Urban MLLM.
@@ 4 Experiments
Figure 4: Qualitative comparisons of generated 3D urban environments from Infinicity, CityGen, CityDreamer and UrbanWorld. By comparison, our method can craft more diverse and realistic 3D urban environments enabling dynamic interactions with humans (walking) and vehicles (driving).
In this section, we first introduce the experimental setup and implementation details (see Section 4.1), and then provide some quantitative evaluations of the created urban environments to demonstrate the superiority of UrbanWorld (see Section 4.2). Finally, we present the generation results of UrbanWorld for qualitative estimation (see Section 4.3).
@@ 4.1 Implementation Details
UrbanWorld incorporates three key techniques: Blender as the professional 3D modeling software, diffusion-based rendering and Urban MLLM-empowered scene design and refinement. Specifically, we use Blender-3.2.2 for Linux systems and the compatible Blosm addon to handle the OSM transformation. In terms of diffusion-based rendering, we utilize Stable Diffusion-1.5 [32] as the fundamental diffusion backbone, combined with ControlNet-Depth [30] when generating multiple views of 3D assets. We also introduce IP-Adapter [33] to support taking reference images as the additional generation condition. We use ControlNet-inpainting [30] as the diffusion controller in the UV texture refinement and ContrlNet-tile [30] in the realness enhancement stage. For the hyper-parameter settings in the rendering part, we set the number of camera views $N=4$, which can basically satisfy the rendering needs of most urban assets. The number of inference steps in all diffusion processes is set as 30 by default. The UV maps of 3D assets are unwrapped in the “smart projection” mode operated in Blender. All experiments are conducted on a single NVIDIA Tesla A100 GPU.
@@ 4.2 Quantitative Evaluations
Method FID (↓) KID (↓) DE (↓) HI (↓) PS (↑)
SGAM [22] 453.81 0.522 0.575 0.872 5.6
PersistentNature [14] 441.65 0.319 0.326 0.742 4.8
SceneDreamer [13] 389.90 0.284 0.152 0.817 6.2
CityDreamer [7] 418.38 0.210 0.147 0.830 6.0
UrbanWorld (text) 377.65 0.187 0.089 0.683 6.5
UrbanWorld (image) 368.72 0.154 0.082 0.665 6.7
To better demonstrate the superior generation performance of UrbanWorld, in this section, we provide quantitative results on five metrics: Frechét Inception Distance (FID), Kernel Inception Distance (KID), Depth error (DE), Homogeneity index (HI) and Preference score (PS). (more details can be found in Appendix A.1).
@@
We test two versions of our method including UrbanWorld (text) which takes textual prompts for generation and UrbanWorld (image) which takes reference images as generation conditions. The compared methods for urban scene generation include SGAM [22], PersistentNature [14], SceneDreamer [8] and Citydreamer [7], representing the most advanced performance of automatic 3D urban environment generation. We don’t provide results from other methods such as CityGen [8], CityCraft [17] and SceneCraft [16] because the source codes are not open up to now.
@@
From the results presented in Table 2, we can observe that UrbanWorld outperforms on each quantitative metric compared with baselines. UrbanWorld with real street-view images as conditions demonstrates improved performance compared to the version relying solely on textual conditions. In detail, compared with the most competitive baseline, UrbanWorld has 5.4% and 26.7% improvement on FID and KID, indicating better realism of generated results. Besides, UrbanWorld achieves 44.2% improvement on depth error, demonstrating the geometry-preserving ability of UrbanWorld. By comparison, rendering methods such as SceneDreamer and CityDreamer can produce visually appealing urban scenes, but commonly lose geometry consistency. In terms of the homogeneity index, UrbanWorld has 10.4% improvement compared with baselines. Consistent with the observation on the qualitative results in Section 4.3, the generated scenes from existing methods exhibit high homogeneity, limited to the style of training data. By comparison, UrbanWorld can produce more diverse urban environments according to user instruction, effectively achieving customized creation. This makes it possible to craft any type of 3D urban environment that adapts to the needs of different urban environments. Lastly, the created urban environments from UrbanWorld obtain higher preference of GPT-4 outperforming all compared methods on the preference score.
@@
Table 3: Study of the effectiveness of three key designs in UrbanWorld with text prompts as generation conditions: Urban MLLM-empowered scene design, texture enhancement and MLLM-assisted scene refinement.
Method FID (↓) KID (↓) DE (↓) HI (↓) PS (↑)
UrbanWorld (text) 377.65 0.187 0.089 0.683 6.5
w/o Urban MLLM design 401.58 0.237 0.104 0.701 6.1
w/o texture enhancement 382.09 0.197 0.125 0.687 6.2
w/o scene refinement 393.73 0.202 0.096 0.690 6.3
Table 4: Study of the effectiveness of three key designs in UrbanWorld with real street-view images as generation conditions: Urban MLLM-empowered scene design, texture enhancement and MLLM-assisted scene refinement.
Method FID (↓) KID (↓) DE (↓) HI (↓) PS (↑)
UrbanWorld (image) 368.72 0.154 0.082 0.665 6.7
w/o Urban MLLM design 373.21 0.158 0.104 0.701 6.5
w/o texture enhancement 384.10 0.172 0.125 0.687 6.5
w/o scene refinement 374.52 0.159 0.096 0.690 6.6
@@
Effectiveness of three key designs. To further validate the effectiveness of key designs in UrbanWorld, we conduct ablation studies and show the results on UrbanWorld (text) and UrbanWorld (image) in Table 3 and 4, respectively. We explore the influence of three designs on the performance, including Urban MLLM-empowered urban scene design, texture enhancement and MLLM-assisted scene refinement. The results indicate that all these techniques contribute to the final generation performance. Specifically, the scene design from Urban MLLM contributes most to the quality of generated urban environments, benefiting from the rich urban environment knowledge of Urban MLLM. The texture completion and enhancement operation have the most notable effect on the estimated depth error because better texture fidelity can help with geometric perception. Besides, the final scene refinement process leads to a gain in all evaluation metrics, further promoting the generation quality.
Figure 5: Illustration of the controllable generation of diverse architecture styles when prompting with reference images (upper left) in UrbanWorld.
Figure 6: Illustration of the controllable generation of diverse urban functional scenes guided by different textual prompts in UrbanWorld.
4.3 Qualitative Results
We present generation results of UrbanWorld produced with textual prompts in Figure 4, including three representative urban functional spaces: residential areas, commercial blocks and parks. For intuitive comparison, we also provide some generation samples from Infinicity [6], CityGen [8] and CityDreamer [7]. The results of Infinicity and CityGen are taken from the original paper because the codes are not open-source. It can be seen that scenes from Infinicity are short of clear textures and well-maintained building structures. Scenes from CityGen feature homogeneous styles without clear characteristics of urban functions. Similarly, the visual appearance of urban elements (especially buildings) in the environments from CityDreamer lacks diversity and is hard to distinguish. Besides, there are also clear geometric distortions of the building boundaries in CityDreamer. These issues will pose great challenges for the real interactions between subjects and urban environments. For example, embodied agents are hard to be trained to conduct urban navigation because the surrounding elements are too similar to recognize. By comparison, the urban elements created by UrbanWorld possess high visual diversity, conveniently controlled by user instructions. Moreover, we explore generating realistic urban assets by prompting UrbanWorld with realistic images as prompts. Figure 5 presents generated urban buildings in various styles guided by realistic images. We can observe that the generated asset possesses high fidelity to the real ones, which is rarely reached by existing approaches. We further extend to generate large-scale urban environments, as illustrated in Figure 6, which presents four representative urban functional spaces, each with a visually consistent and contextually appropriate appearance. These results highlight the superior controllability of UrbanWorld which supports flexible prompts as generation conditions.
4.4 Interactivity Demonstration
Figure 7: Illustration of the interactive nature of created environments from UrbanWorld, showcasing two representative types of interactions: agent perception and navigation.
To demonstrate the interactive characteristics of 3D urban environments produced by UrbanWorld, we provide a case study about how UrbanWorld provides information feedback and embodied activity spaces for the agents. We focus on two representative interaction modes between agents and urban environments: perception and navigation, following existing works about urban agents [15, 2]. As shown in Figure 8, the created urban environment can provide multimodal observations for agents including RGB imagery, depth maps and semantic maps. Such information can be utilized to enhance the perception ability of agents in various complex urban environments, benefiting in accomplishing various embodied tasks such as navigation and object manipulation. We also evaluate the agent’s navigation capabilities within the generated urban environment, where the task is to navigate to specified target coordinates in 3D environments. Here we utilize the Rapidly-exploring Random Tree (RRT) algorithm for path planning using the start coordinates of agents, target coordinates and obstacle coordinates in 3D space. We record the visual observation trajectory of the agent during navigation as shown in Figure 8, demonstrating the effective interaction between agents and environments.
@@ 5 Conclusion
We propose UrbanWorld, the first generative urban world model to create realistic, customized and interactive 3D urban environments with flexible control conditions in a fully automatic manner. Integrating the powerful urban environment understanding ability of an urban MLLM and the controllable generation ability of diffusion models, UrbanWorld can effectively craft high-fidelity and customized urban environments outperforming existing 3D city generation methods. For practical usage, the created urban environments can provide high-fidelity data and interactive environments for developing embodied intelligence and AI agents. We have contributed UrbanWorld as an open-source tool to benefit broad research communities, which we believe can pave a new way to efficiently establish 3D urban world, accelerating the development of AGI.
Appendix A Appendix
A.1 Details of the metrics for quantitative evaluation
Frechét Inception Distance (FID) and Kernel Inception Distance (KID). Both metrics quantify the similarity between the distribution of generated images and real images, where lower values indicate better image quality of generated results. We compute FID and KID between frames sampled from the generated scene and an evaluation set comprising 1000 real street-view images randomly sampled from Google Maps. These metrics can effectively measure the realism of generated results.
Depth error (DE). Depth error is utilized to evaluate the 3D geometry accuracy, following the implementation of EG3D [34] and CityDreamer [7]. Specifically, we use the pre-trained depth estimation model [35] to obtain the “ground truth” of depth maps via density accumulation. DE is then calculated as the L2 distance between the normalized predicted depth maps and the “ground truth”. The final result is averaged on the result from 100 captured frames of generated urban scenes.
Homogeneity index (HI). Realistic cities are featured by complex elements with diverse visual appearances, indicating various functional uses of different urban areas. In order to capture this key character, we propose to evaluate the homogeneity of generated scenes, mainly measuring the variance of different urban scenes. To be specific, we first extract the visual feature of each generated scene image with ResNet [36]. The homogeneity index is then calculated as the averaged cosine similarity of each pair of scenes in the feature space. The smaller value of the homogeneity index means a higher diversity of generated urban environments.
Preference score (PS). Another widely used approach to evaluate the generated result is utilizing powerful LLMs as the evaluator [37, 38], here we prompt GPT-4 to score the snapshots taken from the generated 3D environment, mainly considering the texture sophistication and geometric completeness (ranging from 1 to 10). A higher score indicates stronger preference for the generated result.
A.2 Designed Prompts for urban scene design using Urban MLLM
Figure 8: Designed prompt template for urban scene design using Urban MLLM.